### v1

# Self-trapping and skin solitons in two-dimensional non-Hermitian lattices

Abstract
In the context of non-Hermitian photonics, we systematically investigate the spectral and dynamical aspects of two-dimensional nonlinear Hatano-Nelson lattices. In particular, we examine the interplay between self-trapping, due to nonlinearity, and propagation due to the non-Hermitian skin effect. These antagonistic effects give rise to amplitude thresholds - above which self-trapping in a single-channel dominates - that depend highly on the position of initial excitation and the degree of coupling asymmetry. Additionally, we identify two-dimensional skin soliton solutions and characterize their power thresholds and spatial profiles.
preprint: APS/123-QED
I INTRODUCTION
In recent years, photonics has emerged as a versatile platform for exploring non-Hermitian physics, enabling the precise engineering of gain and loss distributions in optical systems [1]. Unlike quantum or condensed matter systems, where the experimental control of non-Hermiticity remains challenging, photonic systems offer a practical and accessible setup [2]. This has fueled rapid progress in the study of non-Hermitian wave dynamics and the demonstration of novel phenomena, such as exceptional points [3, 4], parity-time $(\mathcal{PT})$ symmetry [5, 6, 7, 8, 9, 10, 11, 12, 13], and the Non-Hermitian Skin Effect (NHSE) [14, 15, 16, 17], all of which are now of crucial importance and relevance to the fields of topological and integrated photonics.
Among the plethora of systems studied in non-Hermitian physics, the Hatano-Nelson (HN) model offers a prototypical lattice, where non-Hermiticity is realized through asymmetric nearest-neighbor couplings [18, 19]. An intriguing feature of this model is the NHSE, i.e., the exponential localization of eigenmodes at the boundary of the system under open boundary conditions (OBC). Initially proposed to describe localization-delocalization transitions in condensed matter systems, the HN model remained for years a purely theoretical concept in mathematical physics [20]. However, recent advances in optics have enabled its experimental realization [21, 22], opening new possibilities for practical applications in photonic cavities and lasers [23, 24, 25, 26], as well as, ultracold atoms [27].
The extension of the NHSE to two-dimensional (2D) systems reveals even richer localization features, that depend on the direction and strength of coupling asymmetry [28, 29, 30]. When the coupling asymmetry is along a single direction, NHSE manifests as side-localized modes; however, when asymmetric couplings are present in both directions, the eigenmodes exhibit significant spatial asymmetry as they become localized at specific lattice corners. Notably, most studies related to higher-dimensional lattices have focused on the topological nature of the NHSE, exploring bulk–boundary correspondence and redefining topological invariants, [31, 32, 33, 34, 35, 36, 37], while much less attention has been devoted to transport features.
On the other hand, nonlinearity, an inherent property of optical systems, gives rise to a wide range of intriguing phenomena, including self-trapping, modulational instability, and soliton formation [38]. When combined with non-Hermiticity, the nonlinear phenomena are significantly modified, leading to novel phenomena and the emergence of localized states [39, 40, 41]. In one-dimensional (1D) HN systems, nonlinearity has primarily been explored in the context of topological edge states and single-mode lasing [42, 43, 44, 45]. Recent studies have also examined the effect of Kerr nonlinearity in 1D HN lattices, focusing on soliton formation [46, 47, 48, 49, 50]. Notably, it was recently shown, both theoretically and experimentally, that the interplay between Kerr nonlinearity and asymmetric couplings can lead to the formation of nonlinear skin solitons [46, 47], characterized by position-dependent power thresholds and highly asymmetric spatial profiles. Nevertheless, the interplay of nonlinearity and non-Hermiticity in higher-dimensional systems, such as 2D HN lattices, remains largely unexplored.
In this work, we study the dynamics in 2D HN lattices with Kerr nonlinearity, focusing on the antagonism between self-trapping, and propagation due to asymmetric couplings. In the extensively studied Hermitian case, where the evolution is described by the discrete nonlinear Schrödinger equation (DNLSE), self-trapping occurs when the amplitude of excitation exceeds a critical threshold, suppressing thus discrete diffraction. This phenomenon has been widely explored in both 1D and 2D systems within the framework of lattice solitons [51, 52, 53, 54, 55, 56, 57]. In particular, we examine how the degree of non-Hermiticity and the location of the initial excitation influence the amplitude thresholds for self-trapping. Our study reveals that these thresholds are highly position-dependent, with relatively low thresholds near the lattice corners toward which the couplings are stronger, and significantly higher thresholds — or even impossible self-trapping — in the bulk of the system, under strong non-Hermiticity. Additionally, we identify two-dimensional skin soliton solutions and characterize their properties through their corresponding power-eigenvalue diagrams. Our results show that these skin solitons exhibit spatial asymmetry, and their power thresholds increase monotonically as the coupling asymmetry becomes stronger. We believe that our study may be highly relevant to the emerging direction of nonlinear and non-Hermitian photonics, where the impact of coupling asymmetry to the lasing skin modes is under investigation.
II TWO-DIMENSIONAL NONLINEAR HATANO-NELSON LATTICE
Figure 1: Schematic representation of a two-dimensional Hatano-Nelson lattice. The arrows represent the asymmetric couplings: $e^{h_{x}}$ and $e^{-h_{x}}$ along the $n_{x}$-direction, and $e^{h_{y}}$ and $e^{-h_{y}}$ along the $n_{y}$-direction. The figure identifies the four geometrically distinct corners of the lattice, labeled as (A), (Bx), (By), and (C). The spatial distribution of four indicative linear eigenmodes is included as inset.
Figure 2: Propagation in a 2D HN lattice for the linear (top row) and nonlinear (bottom row) cases, with non-Hermiticity parameter $h=0.2$, and initial condition $\psi_{n_{x},n_{y}}(z=0)=A\delta_{n_{x},13}\delta_{n_{y},13}$. (a)-(c) Normalized complex amplitude $|\psi_{n_{x},n_{y}}(z)|$ for the linear case at propagation distances (a) $z=1$, (b) $z=3.5$, and (c) $z=6$. (d)-(f) Normalized complex amplitude $|\psi_{n_{x},n_{y}}(z)|$ for the nonlinear case and input amplitude $A=4$, for the same propagation distances as before.
II.1 Coupled mode equations
We begin our study with a two-dimensional generalized nonlinear Hatano-Nelson lattice (2D NLHN) consisting of $N\times N$ coupled waveguides, indexed by $(n_{x},n_{y})\in\{1,2,\dots,N\}$, and characterized by asymmetric couplings. The discrete paraxial equation describing the wave evolution in this lattice, in normalized units, is
$\begin{split}i\frac{\partial\psi_{n_{x},n_{y}}}{\partial z}+e^{-h_{x}}\psi_{n_{x}+1,n_{y}}+e^{h_{x}}\psi_{n_{x}-1,n_{y}}+e^{-h_{y}}\psi_{n_{x},n_{y}+1}\\ +e^{h_{x}}\psi_{n_{x},n_{y}-1}+g|\psi_{n_{x},n_{y}}|^{2}\psi_{n_{x},n_{y}}=0\end{split}$ (1)
where $h_{x},h_{y}\in\mathbb{R}^{+}$ are the non-Hermiticity parameters, in the horizontal ($n_{x}$) and vertical ($n_{y}$) directions, respectively. The quantity $\psi_{n_{x},n_{y}}(z)$ denotes the complex amplitude of the electric field envelope at site $(n_{x},n_{y})$ at propagation distance $z$. The parameter $g$ determines whether the lattice is linear ($g=0$) or exhibits Kerr-type nonlinearity ($g=1$). Throughout this work, we consistently consider OBC in both lattice directions, i.e., $\psi_{0,n_{y}}=\psi_{N+1,n_{y}}=\psi_{n_{x},0}=\psi_{n_{x},N+1}=0$.
The geometry of such a square lattice, schematically depicted in Fig. 1, implies the existence of four geometrically distinct corners. These corners are labeled as follows: corner (A), located in the direction of stronger coupling along both $n_{x}$ and $n_{y}$, where the linear skin modes are localized; corner (C), located in the direction of weaker coupling along both $n_{x}$ and $n_{y}$; and corners (Bx)/(By), located in the directions of stronger coupling along $n_{x}$/$n_{y}$ but weaker coupling along $n_{y}$/$n_{x}$. In the special case where $h_{x}=h_{y}$, corners Bx and By are equivalent by symmetry, and we refer to them in such case as corners (B). We consistently follow this notation throughout this work.
At this point it is important to mention, that this model can be mapped via the imaginary gauge transformation $\alpha_{n_{x},n_{y}}\equiv\psi_{n_{x},n_{y}}e^{-(h_{x}n_{x}+h_{y}n_{y})}$, to a Hermitian tight-binding model, which, in the nonlinear case ($g=1$), is characterized by site-dependent nonlinearity. The complex amplitude evolution in such a lattice, is governed by the following discrete nonlinear Schrödinger (DNLSE) type of equation:
$\begin{split}i\frac{\partial\alpha_{n_{x},n_{y}}}{\partial z}+\alpha_{n_{x}+1,n_{y}}+\alpha_{n_{x}-1,n_{y}}+\alpha_{n_{x},n_{y}+1}\\ +\alpha_{n_{x},n_{y}-1}+ge^{2(h_{x}n_{x}+h_{y}n_{y})}|\alpha_{n_{x},n_{y}}|^{2}\alpha_{n_{x},n_{y}}=0.\end{split}$ (2)
II.2 Impact of nonlinearity on propagation dynamics
In this section, we investigate how the Kerr nonlinearity affects the dynamics in the 2D NLHN lattice. In all numerical results presented in this study, we consider a lattice consisting of $N\times N=25\times 25$ waveguides and equal non-Hermiticity parameters for the two directions, i.e., $h_{x}=h_{y}\equiv h$.
First, as shown in the top row of Fig. 2, in the absence of Kerr nonlinearity ($g=0$), the NHSE dictates the dynamics. Specifically, for a non-Hermiticity parameter $h=0.2$ and an initial condition $\psi_{n_{x},n_{y}}(z=0)=A\,\delta_{n_{x},13}\delta_{n_{y},13},$ where $A$ is the excitation amplitude, the wave function propagates towards the corner (A) of the lattice, as is physically expected. In the linear regime, the diffraction is, by default, independent of the excitation amplitude $A$, unlike the nonlinear case, where its value plays a critical role. In particular, as illustrated in the bottom row of Fig. 2, for the same initial condition with amplitude $A=4$, the wavefunction remains self-trapped at its initial position for short propagation distances ($z<2$). Then, for $z\gtrsim 2$, the wavefunction splits into two parts: one remains localized at the site of the initial condition, while the other propagates toward corner (A). This behavior clearly indicates an antagonism between self-trapping caused by nonlinearity, and the NHSE, caused by the asymmetric couplings.
Figure 3: Evolution of mean positions and uncertainties in a 2D HN lattice with non-Hermiticity parameter $h=0.2$, for the initial condition $\psi_{n_{x},n_{y}}(z=0)=A\delta_{n_{x},13}\delta_{n_{y},13}$. (a) Mean positions $\langle n_{x}\rangle$ and $\langle n_{y}\rangle$ and (b) position uncertainties $\Delta n_{x}$ and $\Delta n_{y}$ for the linear case, blue line) and nonlinear cases with excitation amplitudes $A=1$ (red line), $A=4$ (yellow line), and $A=6$ (magenta line).
Figure 4: Role of $h$ on self-trapping, for fixed position of the excitation channel at the center of the lattice, $\psi_{n_{x},n_{y}}(z=0)=A\delta_{n_{x},13}\delta_{n_{y},13}$. Normalized amplitude $|\psi_{n_{x},n_{y}}(z_{\text{max}})|$ for different values of $A$ (left panels) and $IPR$ versus $A$ (right panels). Red dots correspond to the values of $A$ in the left panels. Two cases are considered: (a) $h=0$ and (b) $h=0.2$.
In order to further understand such an effect, we also study the evolution of the wavefunction’s mean positions and uncertainties along the $n_{x}$ and $n_{y}$ directions. These are defined by the following general expressions:
$\langle n_{x/y}\rangle\equiv\frac{\sum_{n_{x}=1}^{N}\sum_{n_{y}=1}^{N}n_{x/y}|\psi_{n_{x},n_{y}}|^{2}}{\sum_{n_{x}=1}^{N}\sum_{n_{y}=1}^{N}|\psi_{n_{x},n_{y}}|^{2}},$ (3)
$\Delta n_{x/y}\equiv\sqrt{\langle n_{x/y}^{2}\rangle-\langle n_{x/y}\rangle^{2}},$ (4)
where
$\langle n_{x/y}^{2}\rangle\equiv\frac{\sum_{n_{x}=1}^{N}\sum_{n_{y}=1}^{N}n_{x/y}^{2}|\psi_{n_{x},n_{y}}|^{2}}{\sum_{n_{x}=1}^{N}\sum_{n_{y}=1}^{N}|\psi_{n_{x},n_{y}}|^{2}}.$ (5)
In the above expressions, the notation $n_{x/y}$ denotes either $n_{x}$ or $n_{y}$.
Our results are presented in Fig.3. In particular, in Fig. 3(a), we show that, in the linear regime, the wavefunction’s mean positions, $\langle n_{x}\rangle=\langle n_{y}\rangle$, increase almost linearly with the propagation distance $z$, after an initial acceleration [36], and finally the wavepacket reaches the corner C. To the contrary, when Kerr nonlinearity is introduced, propagation is significantly hindered. Specifically, as the excitation amplitude $A$ increases, the mean positions $\langle n_{x}\rangle=\langle n_{y}\rangle$ have a progressively delayed increase and their value never reaches the corresponding for the linear regime, indicating a higher tendency for self-trapping. Moreover, the nonlinear regime is characterized by increased position uncertainties $\Delta n_{x}=\Delta n_{y}$, as depicted in Fig. 3(b). This increase reflects the previously discussed splitting of the wavefunction into two parts (bottom row of Fig. 2). Such behavior emerges due to the interplay between Kerr nonlinearity and the NHSE, which will be studied more systematically in the following section.
III ANTAGONISM BETWEEN SKIN EFFECT AND SELF-TRAPPING
Figure 5: Role of the location of the excitation channel on self-trapping, for fixed $h=0.4$. Amplitude profiles $|\psi_{n_{x},n_{y}}(z_{\text{max}})|$ for various values of $A$ (top-row) under single-channel excitation at different lattice sites (a) at the center, $(n_{x_{0}},n_{y_{0}})=(13,13)$; (b) near corner (A), $(n_{x_{0}},n_{y_{0}})=(23,23)$; and (c) near corner (B), $(n_{x_{0}},n_{y_{0}})=(23,3)$. $IPR$ versus $A$ is shown in bottom-row for all cases. Red dots indicate the values of $IPR$ for the corresponding values of $A$ in top row.
As it was discussed in the previous paragraphs, the skin effect and nonlinearity have opposite tendencies regarding localization, which leads to a dynamic interplay. Therefore, our goal in this section is to demonstrate, through pertinent examples, that the dynamics in a 2D NLHN lattice, as described by Eq. (1) with $g=1$, depend highly on the position of the initial excitation, due to the system’s non-Hermiticity. Specifically, we will consider the single-site initial condition:
$\psi_{n_{x},n_{y}}(z=0)=A\delta_{n_{x},n_{x_{0}}}\delta_{n_{y},n_{y_{0}}}$ (6)
for different cases of $n_{x_{0}}$ and $n_{y_{0}}$ and for various values of the excitation amplitude $A$.
First, we focus on the case of initial excitation at the center of the lattice, i.e., $(n_{x_{0}},n_{y_{0}})=(13,13)$ for Hermitian and non-Hermitian lattices and our results are shown in Fig. 4. In particular, in Fig. 4(a) for the Hermitian case ($h=0$), the wavefunction at a fixed propagation distance $z_{\text{max}}=5$, $|\psi_{n_{x},n_{y}}(z_{\text{max}})|$, is extended for low values of the excitation amplitude $A$. When $A$ exceeds a particular value, the effects of Kerr nonlinearity become significant, leading to almost total self-trapping of the wavefunction at the position of initial excitation. The transition between delocalization and self-trapping can be quantified by the increase in the output (i.e., at $z=z_{\text{max}}$) value of the inverse participation ratio ($IPR$). This metric is widely used to indicate localization and, for a given two-dimensional wavefunction $\psi_{n_{x},n_{y}}$, is defined as:
$IPR\equiv\frac{\sum_{n_{x}=1}^{N}\sum_{n_{y}=1}^{N}|\psi_{n_{x},n_{y}}|^{4}}{\left(\sum_{n_{x}=1}^{N}\sum_{n_{y}=1}^{N}|\psi_{n_{x},n_{y}}|^{2}\right)^{2}},$ (7)
ranging from $1/N^{2}$ for a fully extended state $\psi_{n_{x},n_{y}}\propto 1/N$, to 1 for a state localized to a single site, $\psi_{n_{x},n_{y}}\propto\delta_{n_{x},m_{x}}\delta_{n_{y},m_{y}}$. Interestingly, this behavior changes dramatically when coupling asymmetry is introduced, through non-zero values of $h$. For the case of $h=0.2$, as shown in Fig. 4(b), it is evident that propagation predominates over self-trapping for a much higher range of excitation amplitudes $A$, compared to the Hermitian lattice.
For higher values of the non-Hermiticity parameter $h$, and different locations of the single-channel excitation, our results are presented in Fig. 5. More specifically, for $h=0.4$, Kerr nonlinearity can no longer induce self-trapping for single-site excitation at the center of the lattice, as shown in Fig. 5(a). Instead, the NHSE dominates, forcing the wavefunction to delocalize toward the corner (A) of the lattice. This behavior persists even as the excitation amplitude $A$ increases, within a reasonable range of physically relevant values, as indicated by the consistently low values of the $IPR$. Nevertheless, as shown in Fig. 5(b), the competition between nonlinearity and coupling asymmetry leads to different propagation characteristics when the initial excitation is near the corner (A) of the lattice, i.e., $\psi_{n_{x},n_{y}}(z=0)=A\delta_{n_{x},23}\delta_{n_{y},23}$. Contrary to the case of Fig. 5(a), self-trapping becomes possible above an amplitude threshold. This behavior can be attributed to the fact that the initial excitation is indeed near corner (A), towards which the couplings are higher for both spatial directions $n_{x}$ and $n_{y}$. At this point, it is interesting to also consider the case of an initial excitation near corner (B), i.e., at $(n_{x},n_{y})=(23,3)$. There, delocalization of wavefunctions is promoted only along the spatial dimension $n_{y}$, since the wavefunction is already on the side of the lattice towards which the coupling along $n_{x}$ directions higher. As shown in Fig. 5(c), although self-trapping is possible, it requires considerably higher excitation amplitude. This represents an intermediate case between those discussed in Fig.  5(a) and Fig.  5(b).
To conclude, these examples highlight the intricate interplay between Kerr nonlinearity, which induces self-trapping, and non-Hermiticity, which promotes wavefunction delocalization. Up to this point, it is clear that the antagonism between those two effects depends on three different factors: the non-Hermiticity parameter $h$, the location of the initial excitation $(n_{x_{0}},n_{y_{0}})$, and its amplitude $A$. A natural and physically relevant question arising from our discussion is how to quantify such amplitude thresholds required to induce self-trapping at a single lattice site, for given non-Hermiticity parameter $h$.
IV Amplitude thresholds for Self-Trapping
In order to investigate this question systematically, we study the dynamics in a 2D NLHN lattice, for single-site excitation at each site $(n_{x_{0}},n_{y_{0}})$, gradually increasing the excitation amplitude $A$. For each value of $A$, we compute the output $IPR$ (at $z_{\text{max}}=5$) of the wavefunction, $|\psi_{n_{x},n_{y}}(z_{\text{max}})|$, similarly with the previous section. We define the threshold for self-trapping as the lowest value of $A$ for which $IPR>0.8$. Such a criterion is meaningful as it signifies near-perfect self-trapping at a single site, and although the threshold choice is arbitrary, it does not compromise the generality of our results since any similar criterion would lead to identical conclusions. Furthermore, to ensure that the high value of $IPR$ arises from self-trapping rather than exponential localization toward the corner (A) of the lattice, we impose an additional constraint: the shift of the mean position of the wavefunction at $z=z_{\text{max}}$, $r\equiv\sqrt{(\langle n_{x}\rangle-n_{x_{0}})^{2}+(\langle n_{y}\rangle-n_{y_{0}})^{2}}$, must satisfy $r<1$. For reference, we mention that in the Hermitian case ($h=0$) the two aforementioned constraints yield an amplitude threshold of $A=3.15$ for all lattice sites $(n_{x_{0}},n_{y_{0}})$.
Figure 6: Amplitude thresholds $A$ (colormap) required for single- channel self-trapping for all possible site position. We consider two values of non-Hermiticity parameter: (a) $h=0.2$ and (b) $h=0.4$. Black squares in (b) indicate lattice sites where single-channel self-trapping is impossible for excitation amplitudes up to $A=20$.
At first, we follow the previously described procedure for a lattice with $h=0.2$, and present our results in Fig. 6(a) as a comprehensive map. As evident from this figure, self-trapping occurs at relatively low excitation amplitudes, around $A\sim 3$, for lattice sites adjacent to corner (A). However, the threshold increases to approximately $A\sim 4$ for lattice sites near corners (B) and rises dramatically to $A\sim 9.5$ for the bulk of the lattice. When the non-Hermiticity parameter is increased to $h=0.4$ (Fig. 6(b)), the previously observed asymmetry in amplitude thresholds becomes significantly more profound. While sites near corner (A) still require an excitation amplitude of $A\sim 3$ for self-trapping, the threshold exceeds $A=10$ for sites near corner (B) and, more interestingly, becomes impossible for the bulk of the lattice. For this latter observation, we examined excitation amplitudes up to $A=20$. At this point we note, that when asymmetric couplings only occurs along one transverse direction, i.e., $h_{x}=0.4$, $h_{y}=0$, then self-trapping becomes feasible for all lattice sites.
In conclusion, we have examined how Kerr nonlinearity induces self-trapping in a single-site when the excitation amplitude $A$ exceeds a specific threshold, depending on the location of the initial condition, in a 2D NLHN lattice. This analysis naturally leads to the question of whether this lattice also supports soliton solutions, analogous to the skin solitons recently predicted theoretically [46] and observed experimentally [47] in 1D NLHN systems.
V TWO-DIMENSIONAL SKIN SOLITONS
Figure 7: Power-eigenvalue $P-\mu$ diagrams for different families of solitons (a) near corner (A) $(n_{x_{0}},n_{y_{0}})=(23,23)$ and (b) in the bulk, for a 2D NLHN lattice. Colored lines correspond to different values of the non-Hermiticity parameter $h$. (c) Power thresholds $P_{th}$ as a function of $h$, for solitons near corner (A) (orange line) and in the bulk (purple line). Red, blue and green dots correspond to the dots of (a)-(b) indicating the value of $h$.
In this section, we identify stationary soliton solutions within the 2D NLHN lattice. In particular, we consider solutions of Eq. (1) ($g=1$) of the form $\psi_{n_{x},n_{y}}(z)=\phi_{n_{x},n_{y}}e^{i\mu z}$, where by $\phi_{n_{x},n_{y}}$ and $\mu$ we denote the soliton profile and the soliton eigenvalue, respectively. Substituting this ansatz into Eq. (1) results in a system of nonlinear algebraic equations,
$\begin{split}e^{h}\phi_{n_{x}-1,n_{y}}+e^{-h}\phi_{n_{x}+1,n_{y}}+e^{h}\phi_{n_{x},n_{y}-1}\\ +e^{-h}\phi_{n_{x},n_{y}+1}+|\phi_{n_{x},n_{y}}|^{2}\phi_{n_{x},n_{y}}=\mu\phi_{n_{x},n_{y}}.\end{split}$ (8)
Using iterative numerical techniques [52, 53, 54, 58], we obtain lattice soliton solutions and thus we can relate their corresponding power ($P\equiv\sum_{n_{x}=1}^{N}\sum_{n_{y}=1}^{N}|\phi_{n_{x},n_{y}}|^{2}$) to the soliton eigenvalue ($\mu$). Such power-eigenvalue diagrams are shown in Fig. 7 for various values of $h$. Specifically, we analyze solitons positioned near corner (A) (Fig. 7(a)) and at the center of the lattice (Fig. 7(b)). In both cases, it is evident that the solitons exhibit power thresholds. Additionally, for both soliton locations, we investigate the dependence of the power thresholds $P_{th}$ on the non-Hermiticity parameter $h$. As illustrated in Fig. 7(c), the power thresholds for solitons located near corner (A) and at the center of the lattice increase monotonically with $h$. Specifically, we confirm that this dependence is parabolic, in good approximation, i.e., $P_{th}\approx ah^{2}+bh+c$. We also identify soliton solutions positioned near corners (B) and (C) of the lattice, having power–eigenvalue diagrams very similar to the solitons at the center of the lattice. For a given value of $h$, the power thresholds for corners (B) and (C) differ by less than $10^{-2}$ from those of the solitons in the center. However, the eigenvalues at these thresholds differ more, reflecting variations in the spatial distributions of the soliton profiles.
Figure 8: Soliton profiles $\big{|}\phi_{n_{x},n_{y}}\big{|}$ for $h=0.4$. Each panel corresponds to a different soliton solution family: (a) near corner (A) $(n_{x_{0}},n_{y_{0}})=(23,23)$, (b) near corner (B) $(n_{x_{0}},n_{y_{0}})=(23,3)$, (c) near corner (C) $(n_{x_{0}},n_{y_{0}})=(3,3)$, and (d) the center of the lattice $(n_{x_{0}},n_{y_{0}})=(13,13)$
Furthermore, in Fig. 8 we demonstrate the soliton profiles with the lowest eigenvalue at high non-Hermiticity parameters for different lattice locations. In all cases, the solitons exhibit spatial asymmetry toward corner (A), where linear modes are localized, due to the existence of asymmetric couplings. We also confirmed that, as one would expect, the degree of asymmetry of the soliton solutions increases when $h$ is increased.
Finally, we investigate the stability of the soliton solutions by computing the deviation parameter defined as $\delta\phi_{n_{x},n_{y}}(z)\equiv\max\left(\big{|}|\phi_{n_{x},n_{y}}(z)|-|\phi_{n_{x},n_{y}}(0)|\big{|}\right)$ at a large propagation distance $z_{\text{max}}=25$. This parameter quantifies the deviation of the evolved soliton profile from its initial distribution. For a non-Hermiticity parameter $h=0.2$, solitons at all lattice locations remain stable, exhibiting very small deviations, with $\delta\phi\sim 10^{-7}$. However, at a higher non-Hermiticity value of $h=0.4$, soliton stability varies significantly with lattice position: solitons near corner (A) remain highly stable, with extremely low deviations ($\sim 10^{-10}$), while deviations increase progressively when moving from corner (B) ($\sim 10^{-7}$) toward the lattice center ($\sim 10^{-4}$), reaching their largest value near corner (C) ($\sim 10^{-3}$).
VI Conclusions and discussion
In this work, we have investigated the generalized nonlinear Hatano-Nelson lattice in two spatial dimensions, emphasizing the intricate interplay between skin effect and self-trapping due to Kerr nonlinearity. While in the linear regime, wavefunctions propagate toward a lattice’s preferred corner due to the NHSE, the presence of nonlinearity significantly affects such a preferential direction. We have demonstrated that the single-channel self-trapping strongly depends on the amplitude of the initial excitation, its position, and the degree of coupling asymmetry described. Specifically, initial excitations near the corner where the linear modes are localized, exhibit self-trapping above relatively low amplitude thresholds. Conversely, deep within the lattice bulk, asymmetric couplings actively oppose self-trapping, leading to delocalization even at higher excitation amplitudes; for very strong non-Hermiticity, self-trapping in these bulk regions is practically impossible. Additionally, we have identified and characterized stationary two-dimensional soliton solutions in various lattice locations by computing their associated power-eigenvalue diagrams. These solutions, which generalize the concept of skin solitons in two dimensions, consistently display spatial asymmetry biased toward the lattice corner which is favored by asymmetric couplings. Interestingly, our study revealed that the soliton power thresholds increase parabolically with the non-Hermiticity parameter $h$. Furthermore, solitons positioned near the corner where linear skin modes are localized, remain stable over large propagation distances, whereas those located away from this corner show higher deviations. In conclusion, this study introduces the concept of skin solitons in higher dimensional lattices with asymmetric coupling, setting the stage for further exploration of soliton formation in topologically non-trivial non-Hermitian lattices, such as those exhibiting higher-order skin effects. We believe that our results may provide useful insights regarding transport in higher-dimensional non-Hermitian lattices with nonlinearity, that are highly relevant to the emerging field of non-Hermitian nonlinear optics.
Acknowledgements.
The authors acknowledge financial support from the European Research Council (ERC) through the Consolidator Grant Agreement No. 101045135 (Beyond Anderson). E.T.K. and I.K. further acknowledge funding from the Stavros Niarchos Foundation (SNF) and the Hellenic Foundation for Research and Innovation (H.F.R.I.) under the 5th Call of the “Science and Society” Action, titled “Always strive for excellence – Theodoros Papazoglou” (Project No. 11496, “PSEUDOTOPPOS”). Computations for this paper were partially conducted on the Metropolis cluster, supported by the Institute of Theoretical and Computational Physics, Department of Physics, University of Crete.